# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 5.3 Feature Replay based CIL
Exemplar replay methods suffer from memory limitation and privacy concerns in real-world applications, and generating exemplars in input space is quite challenging and highly depends on the quality of generative models. To avoid directly storing and replaying old data, recent works explore feature replay strategies that store and replay features of old classes. Compared with exemplar replay, feature replay has advantages in terms of computation efficiency, memory footprint and privacy safety. However, with the continual updating of feature extractor in the incremental learning process, the saved old features would be less valid and representative. Therefore, maintaining the effectiveness of saved features is a challenge.
@@
Real and Generative Feature Replay. In DNNs, there exist many features with different levels of depth or semantics. Therefore, both feature instances in intermediate [165, 166] or final layer [167] can be saved. For feature replay in the intermediate layer, to keep the stored features valid and stable when learning new classes, one can directly freeze the layers between input and feature replay layer [166], or assign a lower learning rate for those layers [165]. For feature replay in the final layer, freezing all layers or assigning a lower learning rate would make the model hard to learn new knowledge. To keep the stored features valid, a feature adaptation module can be constructed to map the saved features to the updated feature space [167], and the classifier can be then jointly optimized over new and old classes in the same space. Besides, inspired by generative exemplar replay, some works [168, 169, 170] explored generative feature replay, which generates pseudo features of old classes for replay when learning new classes. In particular, generating the final features has been shown to be more effective than generating features in shallower layers [170].
Prototype Replay. To further reduce the memory cost of feature replay, prototype replay methods [120, 171, 172] only memorize one class-representative prototype (class mean in deep feature space) for each old class. However, the saved prototypes become increasingly outdated when updating the model on new classes. To address this issue, Zhu et al. [120] proposed prototype augmentation to regularize the model, and leveraged self-supervised learning to learn more generic and transferable features, which in turn reduce the feature drift when learning new classes. Other works explicitly model and constrain the feature drifts of old classes via linear [172] or nonlinear transformations [173]. In addition to saving only the prototypes of old classes, more distribution information can be beneficial for maintaining the decision boundary of previous tasks, and an infinite number of pseudo-feature instances can be generated implicitly [174]. More recent works [175, 176] explore directly learning discriminative prototypes for incremental learning.
5.4 Prompt-based CIL
Recent works [177, 178, 179, 180] leverage pre-trained frozen models and learn a set of prompts (i.e., small learnable parameters) to dynamically instruct models in tackling tasks sequentially. L2P [177] stands as the pioneering work that formulates the problem of learning new tasks as training small prompt parameters attached to a pre-trained frozen model. To further improve the plasticity of the model, CODA-Prompt [179] introduces an attention-based end-to-end key-query scheme, wherein a collection of input-conditioned prompts is learned. Wang et al. [181] proposed a hierarchical decomposition prompt framework to explicitly optimize the within-task prediction, task-identity inference, and task-adaptive prediction with task-specific prompts. Nevertheless, it is worth noting that prompt-based methods are largely reliant on pre-trained models. As highlighted recently by Kim et al. [182] the information leak both in terms of features and class labels can exert a substantial influence on the results of those approaches.
6 Future Directions
@@
Open-world machine learning is an active and long-term research topic, and there are a number of critical open directions worthy of further investigation. In this section, we briefly outline a few promising research directions that enable OWL in a unified framework and more complex scenarios, e.g. structured data and applications such as detection, segmentation, etc. Besides, additional directions considering brain-inspired OWL and machine unlearning are also discussed. Figure 6 presents an illustration of the future directions.
6.1 Unified OWL
A desirable open-world learner has to learn and adapt continuously in its interaction with the environment. Specifically, during deployment, the model should correctly classify the known (classes seen at the current stage) and reject the unknown (misclassified samples from known classes and OOD samples from unseen classes), discover and label these unknowns in the buffer as novel known classes, and then incremental learning these new classes while maintaining the discrimination ability on old classes. Nevertheless, existing studies mostly focus on an isolated task without considering the whole learning process. Here we discuss some typical examples. For unknown rejection, current approaches aim to detect and filter out known samples from unseen classes, while they often lower a model’s ability to reject misclassified samples from known classes [58, 183, 184]. This is undesirable because real-world applications require the rejection of both widely existing misclassified and OOD samples, and there is a long way to go towards developing unified unknown rejection methods [185]. For class-incremental learning, when learning new classes, the training samples are well-labeled manually and clean, i.e., only contain given class in $\mathcal{C}^{t}$. In practice, the users often are not aware of novelty and the model is supposed to discover novel classes from the buffer which contains various samples without labels. Some works [186, 187] propose the task of incremental novel category discovery, where the samples from new classes are unlabeled at each incremental step. More recent works [188] consider a more challenging scenario where the unlabeled data could also contain samples from the old classes at each incremental step. In addition, a better understanding of the relationship between unknown rejection, novel class discovery and class-incremental learning is needed. A recent study [182] has shown that OOD detection techniques can help CIL, and designs an algorithm that explicitly performs OOD detection in the incremental learning process. A more interesting question is how rejection performance change when a model continually adapts and learn in the open environment. Intuitively, incremental learning could profoundly affect the rejection ability because newly accumulated knowledge will affect the border of the unknown [12]. In summary, it is worth studying the relation or interaction among those three tasks in OWL and developing a unified framework that can deal with those open-world problems simultaneously, i.e., correctly recognize known, reject unknown, discover new classes from unknown and then incrementally learn them.
@@ 6.1 Unified OWL
Figure 6: Illustrations of future research directions of open-world learning. (A) Unified open-world learner. (B) OWL for foundation models such as GPT, SAM and LLaMA. (C) OWL for structured data with graph neural networks. (D) OWL beyond classification, such as object detection and segmentation. (E) Brain-inspired OWL models that mimic biological learning systems. (F) Open-world unlearning actively and selectively forgetting or erasing specific knowledge.
6.2 OWL for Foundation Models
Recent years have witnessed the significant progress of foundation models [189]. CLIP [190], ChatGPT [23], and SAM [191] have demonstrated remarkable zero-shot and few-shot learning. However, if we consider open-world learning ability, there still remain important limitations. On the one hand, OpenAI has pointed out that ChatGPT can be easily overconfident in its incorrect predictions [23]. When faced with questions it cannot answer, ChatGPT tends to make meaningless guesses rather than refusing to respond [192], propagates false information or provides unethical advice in fields such as clinical and legal matters [24]. Indeed, there is still a lack of clear understanding of the reliability and failure cases of foundational models. On the other hand, foundational models yield poor performance in certain tasks [190, 193]. Nevertheless, they still lack the ability for incremental learning, and direct fine-tuning may deteriorate the model’s performance on other classes or tasks [194]. Furthermore, without the continual updating of model weights, the knowledge within foundational models can become outdated, resulting in the provision of incorrect information. For example, ChatGPT is trained on data up to 2021 and has limited knowledge of the world and events after that time. If you ask ChatGPT who won the World Cup in 2022, it won’t be able to give you a response. Another example is that although ChatGPT is good at writing code, it can not generate code with the latest version, which largely limits its power. Unfortunately, the computational burden of training from scratch makes it impossible to directly retrain the model on both old and new data. The aforementioned issues highlight the limitations associated with the practical applications of foundation models, making it a grand challenge for future research. To enhance the reliability and adaption ability of foundation models in open environments, innovations in both learning objectives and model architectures are needed to develop novel OWL techniques.
6.3 OWL for Structured Data
@@
Most existing OWL methods assume that inputs (e.g. images in computer vision) are independently sampled. However, in practice, data is often structured and inter-connected, such as traffic networks and social networks. However, existing methods are hard to apply to structured data because they mainly focus on individual samples but neglect the topological structure and the interconnections among different samples. There have been some studies that attempt to improve the OWL ability for structured data learning with graph neural networks. To name a few, Stadler et al. [195] designed Bayesian graph neural networks models to detect OOD nodes on graphs, while Wu et al. [196] proposed energy function based OOD discriminator for unknown reject on graph-structured data. For CIL, Liu et al. [197] introduced a topology-aware weight-preserving strategy that explicitly looks into topological aggregation and propagation to reduce the forgetting of old knowledge. Besides, a comprehensive continual graph learning benchmark that contains both node-level and graph-level continual learning tasks has been established recently [198]. Open-world learning on structured data is an emerging research area, and inevitably work will be needed to facilitate its development based on a deeper understanding of structured data.
6.4 OWL Beyond Classification
@@
Existing works mainly focused on image classification tasks. Recently, there has been noticeable attention in addressing open-wold learning challenges in other applications such as object detection [199, 200] and semantic segmentation [201, 202]. In image classification, a single image typically only contains one object. In contrast, object detection and semantic segmentation tasks typically involve input images that contain multiple objects, and some of them are labeled during the initial learning stage, while others should be classified as unknown. This implies that many unknown classes would have been already seen by the object detector. During the incremental learning process, information (labels) about these new classes within the background becomes available, the model should incrementally learn them to incorporate novel knowledge [203]. Therefore, the advances in open-world classification cannot be trivially adapted to open-world object detection or segmentation, and it is important to leverage background information and explore feature representations that could facilitate future learning of new classes. Besides, OWL is also important for natural language processing domain, where out-of-vocabulary words or novel concept vocabulary may emerge over time. To design effective algorithms for those tasks beyond classification, it is helpful to combine existing methods with task-specific domain knowledge.
6.5 Brain-inspired OWL

### v3

@@ V-D Feature Replay Methods
Data replay methods suffer from memory limitations and privacy concerns in real-world applications, and generating exemplars in input space is quite challenging and highly dependent on the quality of generative models. To avoid directly storing and replaying old training data, feature replay strategies have been explored recently. Among them, some methods store and replay feature embeddings in the deep feature space, others only store the prototype, i.e., the class mean of each old class. Moreover, some works developed a generative feature replay technique without strong real feature embeddings. Compared with raw data replay, feature replay has advantages in terms of computation efficiency, memory footprint, and privacy safety. However, with the continual updating of the feature extractor in the continual learning process, the saved old features would be less valid and representative. Therefore, maintaining the effectiveness of saved features is a key challenge for feature replay methods.
@@
1) Real and Generative Feature Replay. In DNNs, there are many features with different levels of depth or semantics. Therefore, both feature instances in intermediate [158, 157] or final layer [160] can be saved. For feature replay in the intermediate layer, to keep the stored features valid and stable when learning new classes, one can directly freeze the layers between input and feature replay layer [157], or assign a lower learning rate for those layers [158]. For feature replay in the final layer, freezing all layers or assigning a lower learning rate would make the model intractable to learn new knowledge. To keep the stored features valid, a feature adaptation module can be constructed to map the saved features to the updated feature space [160], and the classifier can then be jointly optimized over new and old classes in the same feature space. Besides, inspired by generative exemplar replay, some works [152, 368, 153] explored generative feature replay, which generates pseudo features of old classes for replay when learning new classes. For example, Xiang et al. [152] employ an adversarial training scheme to produce mid‑level convolutional features whose distribution closely matches that of old classes. Liu et al. [153] systematically show that synthesizing embeddings in early network layers incurs severe performance degradation, while generating the final features has been shown to be more effective. More recently, van de Ven et al., [159] introduce a brain‑inspired feature replay mechanism that uses the classifier itself to recreate embeddings of old classes, demonstrating strong performance.
Figure 12: Illustration of feature replay methods. (A) Real or generative feature instances replay. (B) Prototype replay.
2) Prototype Replay. To further reduce the memory cost of feature replay, prototype replay methods [162, 164, 161] only memorize one class-representative prototype (class mean in deep feature space) for each old class. Prototype Augmentation and Self‑Supervision (PASS) memorizes a single class‑representative prototype (the class mean) for each old class and applies simple Gaussian‑noise‑based augmentation to preserve the decision boundaries of previous tasks. Specifically, for each old class $k$, pseudo feature instances are generated by augmenting the corresponding prototype:
$\widetilde{\bm{z}}_{k}=\bm{\mu}_{k}+r\cdot\bm{e},$ (25)
where $\bm{e}\sim\mathcal{N}(\bm{0},1)$ is Gaussian noise with the same dimensionality as the prototype $\bm{\mu}_{k}=\frac{1}{n_{k}}\sum\nolimits_{j=1}^{n_{k}}f_{\bm{\theta}}(\bm{x}_{j})$. Hyperparameter $r$ is a scaling factor that controls the level of uncertainty of the augmented features. A key point in the prototype replay is that the saved prototypes become increasingly outdated when updating the model on new classes. To keep stored prototypes valid and representations stable, PASS [162] incorporates self‑supervised learning to obtain more generalizable and transferable features, which in turn reduces feature drift when learning new classes. Other works explicitly model and constrain the feature drifts of old classes via linear [161] or nonlinear transformations [166]. For instance, Toldo et al., [166] model both semantic drift, by learning relationships between novel and existing prototypes, and feature drift, by tracking the evolution of new data embeddings. They jointly exploit these drift estimates to update saved prototypes. In addition to saving only the prototypes of old classes, more distribution information can be beneficial for maintaining the decision boundary of previous tasks, and an infinite number of pseudo-feature instances can be generated implicitly [165]. Some works [68, 92, 180, 369] directly explore learning discriminative prototypes for continual learning. Recent prototype replay methods freeze the backbone and learn a prototype-based classifier. For instance, FeCAM [168] leverages a frozen feature extractor and anisotropic Mahalanobis distance to address heterogeneous feature distributions, while FeTrIL [167] builds a LinearSVCs classifier [370] on the frozen backbone. RanPAC [177] and ACIL [371] transform prototypes of old classes with random projection and then construct a linear regression classifier for all learned classes. Such methods can also be applied to pre-trained models. In general, prototype replay is simple and memory efficient, yet still effective for forgetting mitigation in continual learning.
V-E Model Expansion
1) Backbone Expansion. In the task continual learning setting, dynamically extending the network architecture has been demonstrated to be quite effective [187, 186]. Specifically, when learning a new task, the old parameters are frozen, and new branches are allocated to learn new tasks. However, those methods are often impractical for CL because the task identity is not available at inference time, and one can not know which group of parameters should be used for inference given an input. Recently, based on exemplar replay, several architecture-based methods have been proposed for CL. Adaptive Aggregation Networks (AANets) [183] explores a novel network architecture to explicitly address the stability and plasticity dilemma in class continual learning. Specifically, two types of residual blocks with different numbers of learnable parameters are designed: a stable block to preserve knowledge of old classes, and a plastic block to acquire new classes. Their output feature maps are aggregated before entering the next residual level. A key step in AANets is learning the aggregation weights using a balanced mix of old and new data. A similar idea has been explored in [372], which designed a dual network architecture that includes a slow learning module for generic representation learning and a fast learning module for task-specific representation learning. The slow system is trained via contrastive learning, while the fast system is trained via supervised learning. Dynamically Expandable Representation (DER) [184] dynamically expands the feature extractor to retain old knowledge while learning new concepts. At each incremental stage, the existing feature extractor is frozen, and a new extractor is introduced for the new classes. However, incrementally creating new models leads to an increase in the number of parameters, while KD can be an effective way to remove redundant parameters [138].
Figure 13: Illustration of model expansion methods. (A) Backbone expansion. (B) Prompt-based expansion. (C) LoRA-based expansion. (D) Parameter-efficient expansion for continual learning of multimodal large language models.
2) Parameter Efficient Expansion. There is a rapidly growing paradigm for parameter-efficient adaptation of pre-trained models to continually emerging classes or tasks. These approaches [189, 191, 190, 373, 374, 375] freeze the backbone network and introduce small, learnable modules, either prompts (typically token embeddings maintained in prompt pools) or low-rank adapters to encode task-specific knowledge while minimizing catastrophic forgetting. For example, prompt-based and LoRA (Low-Rank Adaptation) [376] techniques isolate task-specific information, yielding better retention of prior knowledge than direct fine-tuning or classical continual learning methods like data replay and weight regularization. Among them, Learning to Prompt (L2P) [189] stands as the pioneering work that formulates the problem of learning new tasks as training small prompt parameters attached to a pre-trained frozen model. To further improve the plasticity of the model, CODA-Prompt [190] introduces an attention-based end-to-end key-query scheme, wherein a collection of input-conditioned prompts is learned. Wang et al. [197] proposed a hierarchical decomposition prompt framework to explicitly optimize the within-task prediction, task-identity inference, and task-adaptive prediction with task-specific prompts. Guo et al., [202] designed a novel LoRA-based, rehearsal-free continual learning method that integrates new and old knowledge via dynamic representation consolidation and decision boundary refinement, achieving state-of-the-art performance.
Recent methods have further advanced the field through subspace orthogonality constraints to reduce task interference [198, 377, 196, 378], dynamic module allocation and rank selection to manage capacity [198, 203, 379], and sophisticated routing or selection mechanisms, for instance and task agnostic adaptation. InfLoRA [196] achieves interference-free adaptation by designing a subspace that prevents new task updates from impacting old tasks. DualLoRA [377] introduces orthogonal and residual LoRA adapters for each task, controlled by a dynamic memory mechanism to reduce forgetting and enhance learning of new classes. Lu et al. [379] adjust LoRA rank during continual learning to balance stability and plasticity. Nevertheless, it is worth noting that prompt-based methods are largely reliant on pre-trained models. As highlighted by Kim et al. [380], the information leakage, both in terms of features and class labels, can exert a substantial influence on the results of those approaches.
V-F Extended Settings
1) Unsupervised Continual Learning. Conventionally, CL abides by a pure supervised setting, i.e., the training data $\mathcal{D}_{t}=\{\bm{x}_{i}^{t},y_{i}^{t}\}^{n_{t}}_{i=1}$ of each continual stage is entirely labeled. However, the emerging data in real scenarios are not always labeled. Recent works [69, 88, 92] begin to study the setting of unlabeled continual learning. In each continual stage, the model needs to cluster unlabeled samples and discover new classes, while mitigating forgetting the learned old classes in previous states. From this perspective, unsupervised CL could also be referred to as the continual learning version of novel class discovery as discussed in Sec. IV-E. Concurrent works FRoST [68] and NCDwF [69] are the first two works. They combine the self-training algorithms [58] in NCD for continual category discovery and the feature replay [162] spirit to resist forgetting. However, the number of continual learning stages is very limited in their settings [68, 69, 56]. Liu et al. [70] extend this task to scenarios involving multiple stages of continual learning. They discovered that by initializing with high-quality feature representations, freezing the feature extractor and training only the classifier could yield unexpectedly exceptional performance. GM [87] further considers outliers in the evolving training data. In the abovementioned works, the training data of each continual stage exclusively comprises samples from new classes, so their setting could be summarized as the continual version of NCD (Sec. IV-B). However, in real-world applications, the evolving data could simultaneously contain samples from both old and new classes. As a result, the continual version of GCD (Sec. IV-C) is more pragmatic. IGCD [88] focuses on the replay-based setting, and proposes to store class-wise support samples from density peaks. These samples could help overcome forgetting old classes when discovering new ones. Kim et al. [89] utilize proxy-anchor learning schemes for feature representations. They take the spirit of divide-and-conquer and explicitly split old and new classes at each continual stage. PromptCCD [91] devises a Gaussian mixture prompt pool to achieve continual category discovery with prompt learning. Ma et al. [92] study a more challenging and realistic task, namely continual generalized category discovery (C-GCD). In C-GCD, the number of stages in continual learning and the number of unlabeled novel classes are relatively large. C-GCD follows a rehearsal-free CL paradigm, in which the model has no access to previously seen samples due to privacy and memory issues. They incorporate hardness-awareness modeling and propose to replay features of more difficult classes with a higher probability. In this way, the forgetting issue is effectively mitigated, especially for those difficult classes that are more prone to confusion with other classes in the feature space. They further devise group-wise entropy regularization to overcome the model’s prediction bias towards learned classes. In addition, there are a few works that explore semi-supervised [180], self-supervised [381] and federated [192] continual learning.
2) Multimodal Large Language Models. Continual learning of multimodal large language models [382, 383, 384, 385, 386] (MLLM) has emerged as an important research direction to enable lifelong adaptation of large vision-language models. Recent methods focus on parameter-efficient adaptations that overcome catastrophic forgetting as new tasks or domains arrive sequentially. Prompt-based approaches learn modality-specific or fused prompts that isolate task knowledge and facilitate instance or task-aware adaptation without modifying the backbone model [199, 220, 219, 221, 222]. For example, Liu et al. [199] establish a multimodal vision-language continual learning benchmark and propose C-CLIP that consists of multimodal low-rank adaptation and contrastive knowledge consolidation to effectively learn new multimodal tasks while preserving the zero-shot performance of the pretrained model. ModalPrompt [220] introduces prototype prompts for each task and fuses information using image-text supervision. HiDe-LLaVA [219] proposes a task-specific expansion and task-general fusion framework for continual instruction tuning in MLLMs. Recently, Zhao et al. [387] introduced a comprehensive benchmark (MLLM-CL) for continual learning of MLLMs, consisting of two setting named domain and ability continual learning. Then, a multimodal routing mechanism is designed to select task-specific expert in complex multimodal scenarios. These multimodal continual learning methods demonstrate superior performance over traditional strategies, while maintaining scalability and robustness to distributional shifts in practical, real-world settings.
TABLE IX: Comparisons of average accuracies of data replay methods (%) on CIFAR100, ImageNet-Sub, and ImageNet-Full.
Algorithm Venue CIFAR-100 ImageNet-Sub ImageNet-Full
$T=5$ $T=10$ $T=25$ $T=5$ $T=10$ $T=25$ $T=5$ $T=10$ $T=25$
DMC[129] WACV 2020 38.20 23.80 43.07 30.30
GD[349] ICCV 2019 56.39 51.30 58.70 57.70
iCaRL[141] CVPR 2017 57.12 52.66 48.22 65.44 59.88 52.97 51.50 46.89 43.14
BiC[142] CVPR 2019 59.36 54.20 50.00 70.07 64.96 57.73 62.65 58.72 53.47
TPCIL[388] ECCV 2020 65.34 63.58 76.27 74.81 64.89 62.88
UCIR[133] CVPR 2019 63.17 60.14 57.54 70.84 68.32 61.44 64.45 61.57 56.56
WA[143] CVPR 2020 61.70 56.37 50.78 71.26 64.99 53.61 56.69 52.35 44.58
PODnet[134] ECCV 2020 64.83 63.19 60.72 75.54 74.33 68.31 66.95 64.13 59.17
+DDE[137] CVPR 2021 65.42 64.12 76.71 75.41 66.42 64.71
+AANets[183] CVPR 2021 66.31 64.31 62.31 76.96 75.58 71.78 67.73 64.85 61.78
+MRDC[389] ICLR 2022 78.08 76.02 72.72 68.91 66.31
+CwD[390] CVPR 2022 67.44 64.64 62.24 76.91 74.34 67.42 58.18 56.01
Mnemonics[391] CVPR 2020 63.34 62.28 60.96 72.58 71.37 69.74 64.54 63.01 61.00
+AANets[183] CVPR 2021 67.59 65.66 63.35 72.91 71.93 70.70 65.23 63.60 61.53
RMM[355] NeurIPS 2021 68.42 67.17 64.56 73.58 72.83 72.30 65.81 64.10 62.23
DER[184] CVPR 2021 72.60 72.45 77.73
SSIL[146] ICCV 2021 63.02 61.52 58.02
AFC[392] CVPR 2022 66.49 64.98 64.06 76.87 75.75 73.34 68.90 67.02
DMIL[393] CVPR 2022 68.01 66.47 77.20 76.76
V-G Theoretical Study of Continual Learning
With substantial advancements in the empirical methods, recent works also explore the theoretical aspect of continual learning. One line of work considers the setting where all tasks share the same global minimizer. Evron et al. [394] investigated fitting an overparameterized linear regression model for a sequence of tasks with varying input distributions, proving an upper bound of $T^{2}\min\left\{1/\sqrt{k},d/k\right\}$ on the forgetting, where $T$ tasks in $d$ dimensions are presented cyclically for $k$ iterations. Further, Evron et al. [395] studied continual learning of separable linear classification tasks with binary labels, deriving upper bounds on the forgetting effect. Ding et al. [396] focused on the CL of linear regression model via SGD and demonstrated that training tasks with larger eigenvalues in their data covariance matrices later in the sequence increase forgetting when data size is sufficiently large, while an appropriate step size mitigates forgetting. Peng et al. [397] introduced the concept of an ideal continual learner and established its correlation with existing regularization based, data replay based and expansion based methods. The role of Mixture-of-Experts in CL for linear regression tasks under overparameterized regimes has been investigated [398].
Another line of work considers a more general setting where there is no shared global minimize among the continually learned tasks. Within the Neural Tangent Kernel framework [399, 400], the generalization error and forgetting phenomena for orthogonal gradient descent methods [120] have been analyzed. The influence of task similarity on forgetting has been respectively explored for knowledge distillation based methods [401, 402] under overparameterized regimes. Lin et al. [403] provided the explicit form of the forgetting and generalization error for CL of linear regressions with an arbitrary number of tasks, and demonstrated how overparameterization, task similarity, and task ordering affect both forgetting and generalization error. For example, the authors showed that generalization error consistently decreases as tasks become more similar while forgetting can even diminish when tasks are less similar. Deng et al. [404] theoretically demonstrated that sequential rehearsal outperforms standard concurrent rehearsal in mitigating catastrophic forgetting and improving generalization when tasks are less similar in continual learning. Additionally, Kim et al. [405] established a link between task and class continual learning, further providing proof of the learnability of class continual learning [380].
Despite significant advancements, existing theoretical investigations suffer from several fundamental limitations. Firstly, nearly all prior works [394, 395, 396, 397, 403] rely on a naive locally i.i.d. assumption, i.e., the data in each task is required to be independent and identically distributed (i.i.d.) to simplify the analysis, which may be difficult to verify or guarantee in non-stationary nature of real-world data streams, specifically when continually learning and interacting with the dynamic, open environments [9]. Secondly, these studies often impose stringent conditions such as persistent excitation to ensure feature diversity or assume standard Gaussian noise, both of which are difficult to validate in practice. Thirdly, most research focuses on linear regression models, ignoring nonlinear cases (e.g., ReLU/sigmoid activations). Besides, the overparameterized regime assumed in some works lacks applicability to large-scale datasets. A recent work [406] presented a comprehensive theoretical analysis of continual learning for general convex models, and established almost sure convergence results under general data conditions. Moreover, the authors provided the convergence rates for forgetting and regret without excitation assumptions, and derived non-asymptotic error bounds for parameter estimation and CL performance for finite datasets. In future research, a number of theoretical problems remain to be explored, e.g., the convergence properties or estimation error bounds of continual learning with no shared global minimizer, and theoretical investigation for continual learning with DNNs.
TABLE X: Comparisons of the average accuracies (%) of non-exemplar based continual learning methods.
Algorithm Venue CIFAR-100 Tiny-ImageNet ImageNet-Sub
$T=5$ $T=10$ $T=20$ $T=5$ $T=10$ $T=20$ $T=10$
LwF-MC [347] ECCV 2016 33.4 26.0 19.7 34.9 21.4 13.7 35.8
LwM [128] CVPR 2019 39.6 30.2 20.5 37.3 20.5 12.6 32.6
MUC [131] ECCV 2020 49.3 36.0 29.0 37.5 26.3 21.6
CCIL [347] ICME 2021 60.8 43.6 38.1 36.7 27.6 16.3 41.1
UCIR-DF [133] CVPR 2019 57.8 48.7
PODNet-DF [134] ECCV 2020 56.9 52.6
ABD [151] ICCV 2021 62.4 59.0 44.6 41.6
R-DFCIL[366] ECCV 2022 64.8 61.7 48.9 47.6
IL2A [165] NeurIPS 2021 66.2 58.2 58.0 47.2 44.7 40.0 58.0
PASS [162] CVPR 2021 63.8 59.9 58.1 49.5 47.2 42.0 62.1
PASS++ [163] TPAMI 2025 69.1 66.5 64.3 54.1 53.1 49.7 71.9
SSRE [164] CVPR 2022 65.9 65.0 61.7 50.4 48.9 48.2 67.7
SDC [161] CVPR 2020 66.2 62.7 59.2 53.3 50.5 48.8 68.6
Fusion [166] CVPR 2022 66.9 64.8 61.5 54.2 52.6 50.2 69.3
FeTrIL [167] WACV 2023 67.6 66.6 63.5 55.4 54.3 53.0 71.9
Eucl-NCM NeurIPS 2023 64.8 64.6 61.5 54.1 53.8 53.6 72.0
FeCAM [168] NeurIPS 2023 70.9 70.8 69.4 59.6 59.4 59.3 78.2
V-H Evaluation of Continual Learning
In continual learning, performance is typically assessed by last accuracy and average accuracy. Given a dataset containing $M$ classes, one can divide all classes into $T$ disjoint tasks for the model to learn sequentially. Each task contains the same number of classes ($M/T$, assuming $M$ is divisible by $T$). However, motivated by practical scenarios where the initial task often contains more classes, many works assign half of the classes to the initial task and evenly divide the remaining half into $T$ subsequent tasks. Each of these tasks, therefore, contains $M/(2T)$ classes, resulting in a total of $T+1$ tasks. For instance, when applied to CIFAR-100, the initial task would include 50 classes, and the remaining 50 classes would be split into 10 tasks of 5 classes each. Then, at the $t$-th continual stage, the last accuracy $a_{t}$ is defined as the top-1 accuracy of all learned classes at the final continual stage, and the average accuracy is computed as $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{i}$. Following the above evaluation protocols [134, 366], Table IX and Table X provide comprehensive results on common benchmark datasets such as CIFAR-100 [311], Tiny-ImageNet [312], and ImageNet [333] under different continual stages. As can be seen, data replay methods (Table IX) typically achieve good performance, where imbalance calibration is important. However, recent developed non-exemplar methods (Table X) like PASS++ [163] and FeCAM [168] can also yield strong continual learning performance.
VI Future Directions
@@
Open-world learning (OWL) is an active and long-term research topic, it holds transformative potential for developing adaptive artificial intelligence systems that can operate in unpredictable scenarios, from autonomous robotics to personalized healthcare. Over the past decade, substantial strides have been made in endowing systems with the ability to manage the uncertainty inherent in open environments, discover novel concept, and continually update their knowledge. Yet, as OWL moves toward broader real-world deployment, there are a number of critical open directions worthy of further investigation. In this section, we briefly outline a few promising research directions that enable OWL in a unified framework and more complex scenarios (illustrated in Fig. 14), e.g. missing structured data and applications such as detection, segmentation, etc. Besides, additional directions considering brain-inspired OWL and machine unlearning are also discussed.
1) Unified Open-world Learning. A robust open-world learner must continuously adapt through ongoing interaction with its environment, seamlessly integrating classification, rejection, discovery, and continual learning within a single pipeline. During deployment, the model should accurately recognize classes it has encountered so far, while simultaneously rejecting both misclassified examples from known categories and out-of-distribution samples representing unseen classes. Instead of treating rejection and adaptation as isolated tasks, the learner should maintain a buffer of rejected instances and analyze them to identify genuine novelties. By assigning provisional labels to these new classes, the model effectively expands its set of known categories and prepares them for subsequent training. Crucially, as the learner incorporates these discoveries into its knowledge base, it must preserve discrimination performance on previously learned classes and avoid catastrophic forgetting. Existing approaches often address one aspect of this process in isolation. For example, continual learning benchmarks typically assume that new training data are clean and manually labeled according to discrete class sets, whereas real-world applications require autonomous discovery and labeling from an unlabeled stream of mixed samples. Moreover, the dynamic interplay among rejection, novelty discovery, and continual adaptation remains poorly understood. Although recent work [380] has shown that OOD detection techniques can enhance continual learning by filtering novel instances, the shifting decision boundaries that result from accumulating knowledge may in turn affect rejection behavior over time. To move toward truly autonomous, scalable, and reliable open-world learners, it is essential to study the interactions among rejection, discovery, and continual adaptation in concert and to develop unified frameworks that can address all three challenges simultaneously.
@@
Figure 14: Illustrations of future research directions. (A) Unified open-world learner. (B) OWL for large models. (C) OWL for structured data with graph neural networks. (D) OWL beyond classification, such as object detection and segmentation. (E) Brain-inspired OWL models that mimic biological learning systems. (F) Open-world unlearning actively and selectively forgetting or erasing specific knowledge.
2) Open-world Learning of Large Models. Foundation models such as LLaMA [407], DeepSeek [408], LLaVA [409], Qwen [410] and the GPT4 [247] have greatly advanced zero‑shot and few‑shot capabilities in both language and multimodal understanding. However, these models exhibit limitations when deployed in truly open‑world environments where inputs may diverge substantially from their training distributions. One significant challenge is hallucination, which refers to the output of plausible but incorrect or fabricated information [411]. Within an open‑world learning framework, the detection of hallucinations corresponds directly to the problem of unknown rejection, requiring models to withhold responses when confidence is low or when input patterns deviate from established distributions [412]. Another challenge derives from the evolving nature of real‑world knowledge. Without mechanisms for ongoing adaptation, even the most advanced large models can rapidly provide outdated or inaccurate information. To this end, continual pretraining and continual instruction tuning is needed. Specifically, Continual pretraining involves periodically updating model parameters using newly available corpora, thereby ensuring that the model reflects the latest advancements in fields such as science, law and software development [413]. Continual instruction tuning refines a model’s capacity to follow updated user directives and adhere to current protocols [219, 220]. For example, a large model that undergoes continual instruction tuning on recent multimedia search queries will interpret novel query formulations more accurately. To develop OWL methods for large models, we must consider the difference between traditional discriminative models and large generative models, which have more complex generation objectives and diverse forms of task. Furthermore, incorporating user feedback loops, where corrected or clarified responses are incorporated into subsequent instruction‑tuning cycles, can enable models to self‑improve over time. With open-world learning, the next generation of large models will be better equipped to deliver robust, adaptive performance in complex, dynamic scenarios.
@@
3) Open-world Learning of Structured Data. Most open-world learning methods assume that inputs are sampled independently. In many real-world scenarios, however, data exhibit rich structure and interdependence, as seen in traffic networks, social graphs, and molecular interaction maps. Existing open-world approaches struggle with such settings because they focus on isolated instances and overlook the topological relationships that bind samples together. There have been some studies that attempt to improve the OWL ability for structured data learning with graph neural networks. Stadler et al. [414] proposed a Bayesian graph neural network architecture capable of detecting out-of-distribution nodes, while Wu et al. [415] developed an energy-based discriminator to reject unknown inputs in graph-structured data. Liu et al. [416] introduced a topology-aware weight-preservation strategy that explicitly models graph aggregation and propagation, reducing catastrophic forgetting of previously learned relationships during continual learning. Zhang et al. [417] further contributed a comprehensive benchmark for continual graph learning, including both node-level and entire-graph tasks. Open-world learning on structured data is an emerging research area, and future work should leverage a deeper understanding of structured data’s topological properties to develop topology-aware open-world learning methods that can robustly discover and adapt to new knowledge in complex, interconnected scenarios.
@@
4) Open-world Learning Beyond Classification. Existing works mainly focused on image classification tasks. Recently, there has been noticeable attention in addressing open-world learning challenges in other applications, such as object detection [418, 419] and semantic segmentation [420, 421, 422]. In image classification, a single image typically only contains one object. In contrast, object detection and semantic segmentation tasks typically involve input images that contain multiple objects, and some of them are labeled during the initial learning stage, while others should be classified as unknown. This implies that many unknown classes would have already been seen by the object detector. During the continual learning process, information (labels) about these new classes within the background becomes available, the model should incrementally learn them to incorporate novel knowledge [423]. Therefore, the advances in open-world classification cannot be trivially adapted to open-world object detection or segmentation, and it is important to leverage background information and explore feature representations that could facilitate the future learning of new classes. Besides, OWL is also important for the natural language processing domain, where out-of-vocabulary words or novel concept vocabulary may emerge over time. To design effective algorithms for those tasks beyond classification, it is helpful to combine existing methods with task-specific domain knowledge.